OpenStudy (anonymous):

On Monday, a clothing store sold 25 pairs of pants. On Tuesday, the store sold 37 pairs of pants. What was the percent of increase in number of pants sold from Monday to Tuesday?

OpenStudy (anonymous):

pleas help will medal and fan

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (triciaal):

find the change the increase then express that as a percent of the original the change is the difference between the old and the new and percent means per 100 same as setting up the ratio change/original = percentage / 100

OpenStudy (snowsurf):

Let’s write small equation. Let x = percent, P = pants sold, and I= increase in pants sold. \[P + Px = I\]\[P(1+x)=I\]\[1+x=\frac{ I }{ P }\]\[x=\frac{ I }{ P }-1\]\[x=\frac{ I -P }{ P }\] Then plug in values P =25 and I =37 \[x = \frac{ 37-25 }{ 25 }\] x=0.48 Just multiply by 100 to turn in percent.